# GW231123: a product of successive mergers from $\sim 10$ stellar-mass black holes

### v1

@@ Abstract
GW231123 is an exceptionally massive binary black hole (BBH) merger with unusually high component spins. Such extreme properties challenge conventional stellar evolution models predicting a black hole mass gap due to pair-instability supernovae. We analyze GW231123 using population-informed priors on BH mass and spin distributions to test possible formation scenarios: first-generation stellar collapse, hierarchical (multi-generation) mergers, and primordial origin. Our analysis strongly prefers scenarios where at least one component is a higher-generation BH. Both components are favored to have high spins, which rules out scenarios in which they are both first-generation (low spin) or primordial (nearly non-spin). We conclude that GW231123 is a hierarchical merger, with components plausibly originate from the successive mergers of $\sim 6$ and $\sim 4$ first-generation BHs, respectively. This suggests that repeated mergers can be frequent and even more massive intermediate-mass black holes may be produced. Thus mechanisms that can efficiently harden the BBHs’ orbits are required, e.g., gas dynamical fraction in the disks of active galactic nucleus.
@@
software: Bilby (Ashton et al., 2019a, version 1.1.4, ascl:1901.011, https://git.ligo.org/lscsoft/bilby/), Dynesty (Speagle, 2020, version 1.0.1, https://github.com/joshspeagle/dynesty), PyMultiNest (Buchner, 2016, version 2.11, ascl:1606.005, https://github.com/JohannesBuchner/PyMultiNest), Precession (Gerosa & Kesden, 2016, version 1.0.3, https://github.com/dgerosa/precession )
{CJK*}
UTF8gbsn
1 Introduction
Very recently, the LIGO-Virgo-KAGRA (LVK) collaboration reported the detection of GW231123 (The LIGO Scientific Collaboration et al., 2025), a binary black hole (BBH) merger with a remarkably large total source-frame mass of $\sim 190$-$265\,M_{\odot}$ and component dimensionless spins of $\chi_{1}\approx 0.9$ and $\chi_{2}\approx 0.8$. Such extreme masses and spins are unprecedented in previous LVK observations (e.g., Abbott et al., 2023a), and they provide a unique opportunity to probe the formation and evolution of intermediate-mass black holes (IMBHs) (Miller & Colbert, 2004; Liang et al., 2017; Abbott et al., 2022).
@@ 1 Introduction
Alternatively, in dynamical formation channels BHs can undergo repeated mergers and substantially grow in mass, as long as the merger remnants are efficiently retained in their host environments (see Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2025a). GW231123 may thus represent a hierarchical merger. However, how massive BHs can grow via hierarchical assembly remains unclear and likely depends on the host environment. For example, in active galactic nucleus (AGN) disk channels the maximum BH mass may be limited by the finite lifetime of the disk (Xue et al., 2025). Additionally, primordial BHs (PBHs) formed in the early Universe (Carr & Hawking, 1974; Khlopov, 2010) could also populate mass ranges that are hard to reach via stellar collapse (Cai et al., 2018; Escrivà et al., 2024).
@@
Previously, the BBH merger GW190521 (Abbott et al., 2020a) attracted significant attention, as it contained at least one component apparently above the lower edge of PIMG. Fishbach & Holz (2020) and Nitz & Capano (2021) proposed that GW190521 might have been a straddling binary, meaning one BH lay below and the other above the PIMG, so that neither BH was born in the disallowed mass range. However, hierarchical merger scenarios were later found to be more natural for GW190521 once spin information was taken into account. In particular, population analyses inferred that at least one component of GW190521 was likely a higher-generation BH based on its spin and mass properties (Wang et al., 2021b; Kimball et al., 2021; Wang et al., 2022; Li et al., 2024b; Antonini et al., 2025b; Li et al., 2025; Antonini et al., 2025a).
@@
The formation history of a GW event is encoded in the observed source parameters. By performing parameter estimation (inferring the component masses, spins, etc.), one can test various astrophysical origin hypotheses. However, the marginal distributions of these parameters often have large uncertainties (Abbott et al., 2019, 2020b, 2021, 2023a), making it challenging to determine the origins of these events from standard analyses alone. Crucially, some source parameters are correlated or degenerate with each other (Pürrer et al., 2016; Tiwari et al., 2018). For example, adopting a particular prior for the spin distribution can influence the inferred mass distribution for the same event. We therefore employ population-informed priors–priors motivated by the astrophysical distribution of BH spins and masses–to better identify the most plausible formation scenario for GW231123. By examining the joint mass-spin posterior under different prior assumptions, we can potentially rule out formation channels that would fail to reproduce the observed joint mass-spin distribution (even if those channels might naively appear consistent with the one-dimensional mass or spin posteriors under default priors).
@@
In this work, we investigate the origins of GW231123’s two components and the formation channel of the system by analyzing its mass and spin properties with population-informed priors. The rest of the paper is organized as follows: In Section 2, we define the origin scenarios and their expected mass-spin signatures, and describe our population-informed prior choices. In Section 3, we present the results of applying these priors to GW231123 and compare the evidence for each scenario. Besides, we also estimate the number of progenitors needed to generate GW231123 in the hierarchical merger scenario. Finally, we discuss the implications and draw conclusions in Section 4.
2 Population Information to Test Formation Hypotheses
In this section, we introduce the population-informed priors used to represent different formation hypotheses for GW231123.
@@ 2 Population Information to Test Formation Hypotheses
The natal spins of BHs are expected to differ based on their formation mechanism. Stellar-collapse (first-generation) BHs are generally predicted to be born with low spins ($\chi\lesssim 0.1$; Fuller & Ma, 2019), although binary interactions can spin them up modestly (Qin et al., 2018; Bavera et al., 2020; Shao & Li, 2022). In contrast, BHs formed by previous BH-BH mergers tend to have significantly larger spins, typically peaking around $\chi\sim 0.7$ (Gerosa & Berti, 2017; Fishbach et al., 2017; Gerosa & Fishbach, 2021). Indeed, recent population studies of GWTC-3 have found evidence for two distinct subpopulations in the plane of spin magnitude versus component mass (Li et al., 2024b, a; Pierra et al., 2024). The first subpopulation (with smaller BH masses and $\chi\lesssim 0.3$) is consistent with BHs born from stellar core-collapse, whereas the second subpopulation (with more massive BHs and $\chi\sim 0.75$) is indicative of BHs that have undergone at least one merger already.
Motivated by these findings, we introduce spin-magnitude priors corresponding to the low-spin (first-generation) and high-spin (higher-generation) BH populations. Specifically, we take $P_{\rm LS}(\chi)\sim\mathcal{G}(\mu=0.15,\sigma=0.2)$ and $P_{\rm HS}(\chi)\sim\mathcal{G}(\mu=0.8,\sigma=0.3)$, where $\mathcal{G}$ denotes a Gaussian distribution (truncated to the physical range $[0,1]$). These choices are consistent with the spin magnitude distributions inferred from the GWTC-3 BBH population (Li et al., 2024b). We use $P_{\rm LS}$ (“low-spin”) to represent spins of BHs from stellar collapse, and $P_{\rm HS}$ (“high-spin”) to represent spins of BHs formed via previous mergers. For the default prior of spin magnitude, we set $P_{\phi}(\chi)\sim\mathcal{U}(0,1)$, aligning with that of (The LIGO Scientific Collaboration et al., 2025), where $\mathcal{U}$ is the uniform distribution.
@@
Figure 1: Spin magnitude vs. component mass distribution of GW231123 compared to those of BBH events in GWTC-3 informed by population model of Li et al. (2024b). The purple and blue stars (shaded areas) denote the mean values ($90\%$ credible intervals) for the primary and secondary components of GW231123. The green shaded regions show the $90\%$ credible regions for component BHs in the GWTC-3 population, and the black (orange) points mark the mean component masses for the primary (secondary) BHs in those events.
Theoretical predictions also suggest that primordial BHs would have been born with effectively negligible spins (Chiba & Yokoyama, 2017; Green & Kavanagh, 2021). To account for a possible primordial-origin scenario, we include an additional spin prior $P_{\rm NS}(\chi)\sim\mathcal{G}(\mu=0,\sigma=0.01)$, representing a nearly zero-spin distribution. (Whether there exists a subpopulation of BBHs with vanishing spins in current catalog remains under debate (Galaudage et al., 2021; Callister et al., 2022; Tong et al., 2022), but we consider this extreme case for completeness.)
Component masses are also critical to determine the origins of the BHs. Stellar-evolution theory predicts a dearth of BHs in the PIMG roughly between $\sim 40$-$60\,M_{\odot}$ and $\sim 120$-$130\,M_{\odot}$ (e.g., Woosley, 2017; Giacobbo et al., 2018; Woosley, 2019; Farmer et al., 2019, 2020). The precise boundaries of this gap are uncertain and sensitive to model details (nuclear reaction rates, stellar rotation, etc.), but its width is expected to be on the order of $80~{}M_{\odot}$ (Farmer et al., 2019). Some GW population analyses suggest the lower edge of the gap could be as low as $\sim 45~{}M_{\odot}$ (Wang et al., 2021b, 2022; Li et al., 2024b; Pierra et al., 2024; Antonini et al., 2025b). Based on these insights, we impose the following mass priors for stellar-formed BHs: $P_{\rm SM}(m)\sim\mathcal{U}(5,45)M_{\odot}$ and $P_{\rm IM}(m)\sim\mathcal{U}(125,300)M_{\odot}$. We use $P_{\rm SM}$ (“stellar-mass”) and $P_{\rm IM}$ (“intermediate-mass”) for BH masses below and above the PIMG, respectively. In this framework, any BH with mass in the range $\sim 45$ $125~{}M_{\odot}$ cannot be produced by ordinary stellar collapse and is presumed to have a non-stellar origin (most likely a hierarchical merger). Different from that of The LIGO Scientific Collaboration et al. (2025), in our inferences, the priors for component masses are uniform distributions in source frame. For the default prior of component mass, we set $P_{\phi}(m)\sim\mathcal{U}(5,300)$.
We note that a purely non-astrophysical origin (both BHs being primordial) (Carr & Hawking, 1974; Khlopov, 2010; Cai et al., 2018; Escrivà et al., 2024) would in principle allow masses in a much broader range than the stellar paradigm. However, because the PBH mass spectrum is not well constrained (De Luca et al., 2021; Yuan et al., 2024), we do not assign a distinct PBH-specific mass prior. Instead, the PBH scenario is tested primarily through the spin prior $P_{\rm NS}(\chi)$, since PBHs would be characterized by nearly zero spins.
Different formation channels for BBHs predict different spin orientation distributions. Isolated field binary evolution (with processes like mass transfer and tidal locking) (see Mandel & Farmer, 2022, and the reference therein) tends to produce BBHs with nearly aligned spins (i.e., small tilt angles between the BH spins and the orbital angular momentum) (Rodriguez et al., 2016). Especially, the chemically homogeneous evolution (CHE) channel, which is potentially able to produce component BHs as massive as GW231123 (Marchant et al., 2016). In contrast, dynamical formation in dense environments (e.g., multi-body interactions in star clusters) typically leads to isotropic spin orientations (random spin tilt angles) (Talbot & Thrane, 2017). An exception is the dynamical formation in AGN disks, where gas torques can align the spins with the orbital axis (Yang et al., 2019; Tagawa et al., 2020a; Li et al., 2022b). In fact, population studies suggest that a fraction of BBH mergers–especially the more massive, hierarchical ones–may occur in AGN environments (Wang et al., 2021a; Gayathri et al., 2021; Wang et al., 2022; Li & Fan, 2025b; Li et al., 2025; Zhu & Chen, 2025).
However, the spin tilts of GW231123 remain poorly constrained: different waveform models yield inconsistent orientation measurements, and independent analyses of LIGO Livingston and Hanford data also disagree (The LIGO Scientific Collaboration et al., 2025). Therefore, we do not implement a population-informed prior for the spin tilt angles for this event.
3 Results
Table 1: Bayes factors of various population-informed priors (x) compared by the default priors ($P_{0}$)
$\ln\mathcal{B}^{P_{0}}_{x}$ $a_{1}\sim P_{\rm NS}(a)$, $a_{1}\sim P_{\rm LS}(a)$, $a_{1}\sim P_{\rm HS}(a)$, $a_{1}\sim P_{\rm LS}(a)$,
$m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\rm IM}(m)$
$a_{2}\sim P_{\rm NS}(a)$,$m_{2}\sim P_{\phi}(m)$ 26.5 15.1 2.8 17.3
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\phi}(m)$ 25.4 12.8 1.7 11.1
$a_{2}\sim P_{\rm HS}(a)$,$m_{2}\sim P_{\phi}(m)$ 5.7 1.5 -0.8 1.8
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm IM}(m)$ 25.0 12.0 1.3 11.2
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm SM}(m)$ 25.4 11.9 1.5 12.3
Note: The Bayes factors with $\ln\mathcal{B}^{P_{0}}_{x}<0$ ($\ln\mathcal{B}^{P_{0}}_{x}<2.3$) are marked in blue (orange), denoting that
the $x$ prior is favored (not ruled out), comparing to the default prior $P_{0}$.
3.1 What is the most plausible origin for GW231123?
Table 1 summarizes the natural logarithm of the Bayes factors for all population-informed priors relative to the default priors. We find that a zero-spin prior is disfavored for either component of GW231123, suggesting that a PBH origin (which would predict non-spinning BHs) is unlikely. Furthermore, a low-spin prior is less favored either. In other words, neither component of GW231123 appears to come from the low-spin subpopulation identified by Li et al. (2024b), which is thought to consist of first-generation (stellar-collapse) BHs.
We find the scenarios that the secondary BH belongs to the low-spin, low-mass subpopulation can not be ruled out. However, the primary component should belong to the high-spin subpopulation. This means that while it is possible for the two components of GW231123 to straddle the PIMG similar to the case of GW190521 (Fishbach & Holz, 2020), the primary BH is not formed directly via the core collapse of two massive stars. Additionally, the case the secondary BH being a low-spin IMBH (consistent with stellar-collapse origin) is also possible, as long as the primary BH is highly spinning (consistent with merger remnant).
@@ 3.1 What is the most plausible origin for GW231123?
All of our results indicate that at least one component of GW231123 belongs to the high-spin subpopulation (identified in GWTC-3, Li et al., 2024b), which is consistent with a second- or higher-generation BH origin. This strongly supports the interpretation that GW231123 is a hierarchical merger. We plot the component-mass versus spin-magnitude distributions of GW231123111The posterior samples are adopted from https://zenodo.org/records/16004263, comparing to that of the previous events reweighed [^2] The data and codes can be download from GitHub: Stellar-formed V.S. Merger-formed by the population model in Li et al. (2024b), see Figure 1. It shows that both components are consistent with the high-spin subpopulation, which strongly supports that both components of GW231123 are higher-generation BHs.
Figure 2: Mass distributions of merger remnants formed from first-generation BHs in dynamical formation channels, compared to the component masses of GW231123. The first-generation BH mass distribution is taken from Li et al. (2024a). The pink and grey shaded regions indicate GW231123’s primary and secondary mass distributions, which align with mass distribution of remnants assembled from $\sim 6$ (green) and $\sim 4$ (blue) first-generation BHs, respectively. A self-consistency check comparing these model predictions to the observed higher-generation BH population is provided in Appendix B (see Figure 4).
3.2 How many stars are needed to generate GW231123?
Theoretical predictions and population analysis suggest that isolated binary evolution channels contribute to the LVK’s BBH mergers (see e.g. Zevin et al., 2021; Wang et al., 2022; Godfrey et al., 2023; Li et al., 2024a), which are not expected to participate in subsequent hierarchical mergers (Gerosa & Fishbach, 2021). Only the BHs in dynamical formation channels may undergo hierarchical mergers, and potentially contribute to the GW231123-like events. Recent population analysis revealed a subpopulation of BHs, which dominates the mass range $\sim[20,40]~{}M_{\odot}$ (see e.g. Wang et al., 2022; Ray et al., 2024; Li et al., 2024a), consistent with dynamical formation channels predicted by simulations (see e.g., Antonini & Gieles, 2020; Antonini et al., 2023). Therefore, in estimating how many smaller BHs are required to build GW231123’s components, we adopt the mass distribution of the BH subpopulation associated with first-generation BHs in dynamical formation channels, as identified by Li et al. (2024a) [^3] Data and codes are available at GitHub: Field V.S. Dyanmical. Figure 2 shows the mass distribution of hierarchical merger remnants produced by BHs drawn from this subpopulation, compared against the observed masses of GW231123. From these distributions, we infer that producing a BH of GW231123’s primary mass would require merging roughly 6 first-generation BHs, while the secondary could be built from about 4 first-generation BHs (implying a total equivalent of $\sim 10$ stellar-mass BHs to assemble the system’s $\sim 250~{}M_{\odot}$ total mass).
For consistency, we also compare these remnant distributions with the subpopulation of higher-generation BHs observed in GWTC-3 (Li et al., 2024b, a). We find that the distribution of higher-generation BH masses is consistent with remnants produced by the mergers with $\gtrsim 2$ first-generation BHs, supporting the idea that many of the heavy BHs in GWTC-3 (and GW231123 itself) are products of hierarchical growth (see Figure 4 in Appendix B for a self-consistency check).
The primary component of GW231123 could also arise from a direct merger between a $\sim 125M_{\odot}$ stellar-collapse (1G) IMBH and a $\sim 35M_{\odot}$ BH, which would shorten the merger chain to generate GW231123. However, the predicted final spin in such a scenario is significantly lower than the measured primary spin of GW231123, see Figure 3 (right panel). Consequently, it is more plausible that both components of GW231123 were assembled through successive mergers of first-generation BHs below the PIMG.

### v2

# GW231123: Likely a product of successive mergers from $\sim 10$ stellar-mass black holes

@@ Abstract
GW231123 is an exceptionally massive binary black hole (BBH) merger with unusually high component spins. Such extreme properties challenge conventional stellar evolution models predicting a black hole mass gap due to pair-instability supernovae. We test possible formation scenarios for GW231123 using population-informed priors on BH spin distributions, in light of population properties built on the previous (GWTC-3) data. Our analysis shows that GW231123 belongs to the high-spin subpopulation that is naturally interpreted as hierarchical BBH mergers. By comparing the spin magnitudes and component masses of GW231123 to those of the remnants of previous mergers, we show that both components of GW231123 are multi-generation (>>2G) merger remnants, and plausibly originated from the successive mergers of $\sim 6$ and $\sim 4$ first-generation BHs, respectively. This suggests that repeated mergers can be frequent and even more massive intermediate-mass black holes may be produced. Thus mechanisms that can efficiently harden the BBHs’ orbits are required, e.g., gas dynamical friction in the disks of active galactic nuclei.
@@
software: Bilby (Ashton et al., 2019, version 1.1.4, ascl:1901.011, https://git.ligo.org/lscsoft/bilby/), Dynesty (Speagle, 2020, version 1.0.1, https://github.com/joshspeagle/dynesty), PyMultiNest (Buchner, 2016, version 2.11, ascl:1606.005, https://github.com/JohannesBuchner/PyMultiNest), Precession (Gerosa & Kesden, 2016, version 1.0.3, https://github.com/dgerosa/precession )
I Introduction
Very recently, the LIGO-Virgo-KAGRA (LVK) collaboration reported the detection of GW231123 (Abac et al., 2025), a binary black hole (BBH) merger with a remarkably large total source-frame mass of $\sim 190$-$265\,M_{\odot}$ and component dimensionless spins of $\chi_{1}\approx 0.9$ and $\chi_{2}\approx 0.8$. Such extreme masses and spins are unprecedented in previous LVK observations (e.g., Abbott et al., 2023a), and they provide a unique opportunity to probe the formation and evolution of intermediate-mass black holes (IMBHs) (Miller & Colbert, 2004; Liang et al., 2017; Abbott et al., 2022).
@@ I Introduction
Alternatively, in dynamical formation channels BHs can undergo repeated mergers and substantially grow in mass, as long as the merger remnants are efficiently retained in their host environments (see Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2025). GW231123 may thus represent a hierarchical merger. However, how massive BHs can grow via hierarchical assembly remains unclear and likely depends on the host environment. For example, in active galactic nucleus (AGN) disk channels the maximum BH mass may be limited by the finite lifetime of the disk (Xue et al., 2025). Additionally, primordial BHs (PBHs) formed in the early Universe (Carr & Hawking, 1974; Khlopov, 2010) could also populate mass ranges that are hard to reach via stellar collapse (Cai et al., 2018; Escrivà et al., 2024).
@@
Previously, the BBH merger GW190521 (Abbott et al., 2020a) attracted significant attention, as it contained at least one component with a mass apparently above the lower edge of the PIMG Fishbach & Holz (2020). Nitz & Capano (2021) proposed that GW190521 might be a straddling binary, meaning one BH lay below and the other above the PIMG. However, population analyses inferred that at least one component of GW190521 was likely a higher-generation BH based on its spin and mass properties (Wang et al., 2021; Kimball et al., 2021; Wang et al., 2022; Li et al., 2024b; Pierra et al., 2024a; Antonini et al., 2025c; Li et al., 2025a; Antonini et al., 2025a).
@@
The formation history of a GW event is encoded in the observed source parameters. By performing parameter estimation (inferring the component masses, spins, etc.), one can test various astrophysical origin hypotheses. However, the marginal distributions of these parameters often have large uncertainties (Abbott et al., 2019, 2020b, 2021, 2023a), making it challenging to determine the origins of these events from standard analyses alone. Crucially, some source parameters are correlated or degenerate with each other (Pürrer et al., 2016; Tiwari et al., 2018). For example, adopting a particular prior for the spin distribution can influence the inferred mass distribution for the same event. We therefore employ population-informed priors – priors in light of population analysis built on the previous BBH catalog – to better understand the origin and evolution history of GW231123.
@@
The rest of the paper is organized as follows: In Section II, we describe our population-informed prior choices, and present the results of applying these priors to GW231123 and compare the evidence for each scenario. Besides, we also estimate the number of progenitors needed to generate GW231123 in the hierarchical merger scenario. In Section III, we discuss the implications and draw conclusions.
II Testing the Formation Hypotheses with Population Information
The spins of BHs are expected to differ based on their formation mechanism. Stellar-collapse (first-generation) BHs are generally predicted to be born with low spins ($\chi\lesssim 0.1$; Fuller & Ma, 2019), although binary interactions can spin them up modestly (Qin et al., 2018; Bavera et al., 2020; Shao & Li, 2022). Primordial BHs would also have effectively negligible spins (Chiba & Yokoyama, 2017; De Luca et al., 2019; Green & Kavanagh, 2021), neglecting any significant accretions that PBHs may undergo (De Luca et al., 2020)  [^1] There is no evidence for an exactly non-spinning subpopulation (Callister et al., 2022; Tong et al., 2022; Hussain et al., 2024) within current GW data. Additionally, it would be challenging to identify mergers of PBHs from astrophysical within era of second-generation interferometers (Franciolini et al., 2022).. In contrast, BHs formed by previous BBH mergers tend to have significantly larger spins, typically peaking around $\chi\sim 0.7$ (Gerosa & Berti, 2017; Fishbach et al., 2017; Gerosa & Fishbach, 2021). Particularly, spin magnitudes of higher-generation BHs assembled in the AGN disks may shift to even larger values due to gas accretion (Vaccaro et al., 2024; Ford & McKernan, 2025; Bartos & Haiman, 2025).
@@ II Testing the Formation Hypotheses with Population Information
Indeed, recent population studies of GWTC-3 have found evidence for two distinct subpopulations in the plane of spin magnitude versus component mass (Li et al., 2024b; Pierra et al., 2024b). The first subpopulation (with smaller BH masses and $\chi\lesssim 0.3$) is consistent with BHs born from stellar core-collapse, whereas the second subpopulation (with more massive BHs and $\chi\sim 0.75$) is indicative of BHs that have undergone at least one merger already. Similar subpopulations are also identified with different approaches (Hussain et al., 2024; Antonini et al., 2025c; Li et al., 2025a), although some population analyses that concentrate on the marginal distributions of spins (or that do not account for the mass-spin correlations) did not find evident subpopulations (e.g. Abbott et al., 2023b; Edelman et al., 2023; Mould et al., 2022; Wysocki et al., 2018; Golomb & Talbot, 2023; Adamcewicz et al., 2024). With the newly released data of GWTC-4, the subpopulations of low-spin and high-spin BHs are more evident (e.g., Tong et al., 2025; Antonini et al., 2025b; Banagiri et al., 2025; Li et al., 2025b; Wang et al., 2025b).
II.1 Both components are likely higher-generation BHs
@@
Figure 1: Spin magnitude v.s. component mass distribution of GW231123 ( reweighed by HS-HS prior) compared to those of BBH events in GWTC-3 informed by population model of Li et al. (2024b). The red and blue contours denote the primary and secondary components of GW231123. The green shaded regions show the $90\%$ credible regions for component BHs in the GWTC-3 population, and the black (orange) points mark the mean component masses for the primary (secondary) BHs in those events.
Motivated by the subpopulations that are identified with previous data (Li et al., 2024b) and the potential links  [^2] Some work also suggested the existence of a high spin population (Roulet et al., 2021; Hoy et al., 2022; Galaudage et al., 2021; Hussain et al., 2024), though some are unlikely hierarchical mergers, due to divergent model assumptions. The identification and interpretation of first-generation and higher-generation BHs is being common (e.g. Wang et al., 2022; Guo et al., 2024; Antonini et al., 2025a), especially with GWTC-4 (e.g., Tong et al., 2025; Antonini et al., 2025b; Afroz & Mukherjee, 2025; Banagiri et al., 2025; Li et al., 2025b; Wang et al., 2025b). to the origins of BHs (Gerosa & Fishbach, 2021), we apply the population-informed priors based on the subpopulations mentioned above to test the origin of GW231123. Specifically, we use the priors informed by the population of BHs inferred from GWTC-3 (Li et al., 2024b), see Appendix B for detailed illustration.
@@
Comparing the Bayes factors between different priors (see Appendix B), we find the most favored scenario is that both BHs of GW231123 belong to the high-spin subpopulation. Considering the origin of GW231123 in light of the subpopulation properties inferred with previous data (Gerosa & Fishbach, 2021; Li et al., 2024b), the interpretation that GW231123 is a hierarchical merger is very natural. Such a conclusion is clearly illustrated in the (population-informed) component-mass versus spin-magnitude distributions of GW231123333The posterior samples are obtained from https://zenodo.org/records/16004263 ligo_scientific_collaboration_2025_16004263 (catalog ), and the ‘Mixed sample’ are adopted., comparing to that of the previous events reweighed [^4] The data and codes can be download from GitHub: Stellar-formed V.S. Merger-formed jacklee0214_2025_17572948 (catalog ), which built on the ‘Mixed samples’ of GWTC-3. by the population model in Li et al. (2024b), see Figure 1.
Under the population-informed priors that both components belong to the high-spin subpopulation (i.e., HS-HS), we obtain $\chi_{1}=0.82^{+0.13}_{-0.22}$ and $\chi_{2}=0.77^{+0.17}_{-0.31}$, see Appendix A for the details of reweighting. We find the component masses are also changed after reweighting, since the source parameters are degenerated in the parameter estimation of individual event. Under HS-HS priors, we find the secondary mass has fewer samples at low-mass range, providing further support for the higher-generation scenario. It is worth mentioning that under the HS-LS prior (assuming the secondary BH is slowly spinning), the secondary mass shows increasing support at $\sim 60M_{\odot}$. However, the HS-LS prior is less favored compared to the HS-HS prior by $\log_{10}\mathcal{B}=-0.7$.
II.2 Likely higher than second-generation BHs
The final spin of the merger remnant also depends on the spin of BH progenitors (Gerosa & Berti, 2017; Fishbach et al., 2017; Mahapatra et al., 2024). Mergers of low-spin equal-mass progenitors have final spin narrowly peaking at $0.7$, while mergers of high-spin progenitors have a broad distribution of final spins. Particularly, if the highly spinning progenitor BBHs tend to align their spins with the orbits, the final spins will shift to larger values. We generate the final spins of the BBHs in three cases based on the subpopulations of Li et al. (2024b): one is that both BHs come from the LS subpopulation (1G+1G), the second is that the two BHs respectively come from LS and HS subpopulations (nG+1G), the third is that both BHs come from the HS subpopulation (nG+nG). We find that both spin magnitudes of GW231123 are more consistent with the remnants of nG+nG or nG+1G rather than the remnants of 1G+1G.